# Principal stratification with $U$-statistics under principal ignorability

### v3

@@ 3.1 Identification of the PGCE estimands
Theorem 1 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al., 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al. (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively. Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Along the lines in Jiang et al. (2022), we define the following quantity for any function $f(Y,D,X)$:
[…]
One example of (10) is having $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis, 2006):
$\widehat{p}_{z}\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})=\mathbb{E}_{n}\left[\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X;\widehat{\beta})\}}{\pi(X;\widehat{\alpha})}+p_{z}(X;\widehat{\beta})\right],$ (12)
which improves upon the plug-in estimator $\mathbb{E}_{n}p_{z}(X;\widehat{\beta})$ in terms of estimation efficiency and robustness, and is consistent for $p_{z}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$.
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998). For example, Theorem 1(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
$\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
Theorem 1(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
and Theorem 1(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
$\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
The constructions of plug-in estimators for the remaining PGCEs are similar and, therefore, omitted. These estimators have stronger requirements on model specification to estimate the PGCEs consistently. Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
Remark 3
@@ 3.2 Estimators based on nonparametric identification formulas
The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al., 2022); see the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
4 Multiply robust and efficient $U$-estimators
Although implementing the plug-in estimators motivated by the nonparametric identification formulas is straightforward, their various forms and potential inefficiency call for alternative estimators with improved statistical properties. Here, we develop alternative $U$-estimators that exploit information across all nuisance functions, including the treatment assignment, the intermediate variable, and the final outcome of interest.
4.1 The efficient influence function for the PGCE
@@
We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis, 2006; Mao, 2017). Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators. Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al., 2018; Rotnitzky et al., 2021). We will explore these properties by first formalizing the EIF for the PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$. We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following results are obtained.
@@
Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
@@
+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
@@
(b) The EIF for $\tau_{h}^{00}$ is $\varphi_{h}^{00}(\mathcal{O})=\{\varphi_{h,N}^{00}(\mathcal{O})-\tau_{h}^{00}\varphi_{D}^{00}(\mathcal{O})\}/(1-p_{1})^{2}$, where $\varphi_{h,N}^{00}(\mathcal{O})=\phi_{h}^{00}(\mathcal{O})-2\tau_{h,N}^{00}$ and
$\phi_{h}^{00}(\mathcal{O})=$ $~{}\frac{Z(1-D)}{\pi(X)}\left\{\nu_{1,00}^{00}(Y)-m_{1}^{00}(X)\right\}+\frac{(1-Z)(1-D)e_{00}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,10}^{00}(Y)-m_{0}^{00}(X)\right\}$
$+\psi_{1-D_{1}}(\mathcal{O})\left\{m_{1}^{00}(X)+m_{0}^{00}(X)\right\},$
with $m_{1}^{00}(X)=\int e_{00}(x)\mu_{1000}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{00}(X)=\int e_{00}(x)\mu_{1000}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{00}(\mathcal{O})=2\psi_{1-D_{1}}(\mathcal{O})(1-p_{1})-2\tau_{D}^{00}.$
(c) The EIF for $\tau_{h}^{11}$ is $\varphi_{h}^{11}(\mathcal{O})=\{\varphi_{h,N}^{11}(\mathcal{O})-\tau_{h}^{11}\varphi_{D}^{11}(\mathcal{O})\}/p_{0}^{2}$, where $\varphi_{h,N}^{11}(\mathcal{O})=\phi_{h}^{11}(\mathcal{O})-2\tau_{h,N}^{11}$ and
$\phi_{h}^{11}(\mathcal{O})=$ $~{}\frac{ZDe_{11}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,01}^{11}(Y)-m_{1}^{11}(X)\right\}+\frac{(1-Z)D}{1-\pi(X)}\left\{\nu_{0,11}^{11}(Y)-m_{0}^{11}(X)\right\}$
$+\psi_{D_{0}}(\mathcal{O})\left\{m_{1}^{11}(X)+m_{0}^{11}(X)\right\},$
with $m_{1}^{11}(X)=\int e_{11}(x)\mu_{1101}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{11}(X)=\int e_{11}(x)\mu_{1101}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
The EIFs in Theorem 2 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al. (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
$\tau_{h}^{s}=\mathbb{E}\{\mathfrak{h}_{1}^{s}(Y)|Z=1\}+\mathbb{E}\{\mathfrak{h}_{0}^{s}(Y)|Z=0\},$
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 2, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4
Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
$\phi_{h}(\mathcal{O})=\frac{Z}{\pi(X)}\left\{\nu_{1}(Y)-m_{1}(X)\right\}+\frac{1-Z}{1-\pi(X)}\left\{\nu_{0}(Y)-m_{0}(X)\right\}+m_{1}(X)+m_{0}(X),$
with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$. The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$. This coincides with the EIF developed in Mao (2017) using the projection approach (Tsiatis, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
@@ 4.2 Multiply robust estimators
The EIFs given in Theorem 2 imply the semiparametric efficiency lower bound and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 2, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}},~{}~{}~{}\tau_{h}^{00}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{00})}{\{\mathbb{E}(\psi_{1-D_{1}})\}^{2}},~{}~{}~{}\tau_{h}^{11}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{11})}{\{\mathbb{E}(\psi_{D_{0}})\}^{2}}.$ (13)
Continuing with parametric models for nuisance functions, the denominators in (13) can be estimated using the doubly robust estimator given in (12). Specifically, let $\theta$ denote the collection of parametric nuisance function parameters and $\vartheta$ denote a subset of $\theta$, i.e., $\theta=\{\alpha,\beta,\gamma\}$ and $\vartheta=\{\alpha,\beta\}$, with corresponding maximum likelihood estimators $\widehat{\theta}=\{\widehat{\alpha},\widehat{\beta},\widehat{\gamma}\}$ and $\widehat{\vartheta}=\{\widehat{\alpha},\widehat{\beta}\}$, as well as true values $\theta^{*}=\{\alpha^{*},\beta^{*},\gamma^{*}\}$ and $\vartheta^{*}=\{\alpha^{*},\beta^{*}\}$. Then, for example, the denominator of $\tau_{h}^{10}$ given in (13) can be estimated by $[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}$, where $\mathbb{E}_{n}\psi_{D_{z}}(\widehat{\vartheta})\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})$ defined in (12).
Estimating the numerators in (13) is more complicated. To target these terms, we construct the following functions of the observed pairs:
@@
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{00}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
$=2^{-1}\bigg{[}\frac{Z_{i}(1-D_{i})}{\pi(X_{i};\alpha)}\times\frac{(1-Z_{j})(1-D_{j})e_{00}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1000}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}(1-D_{j})}{\pi(X_{j};\alpha)}\times\frac{(1-Z_{i})(1-D_{i})e_{00}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1000}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1000}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1000}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{11}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ $=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{11}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})D_{j}}{1-\pi(X_{j};\alpha)}\left\{h(Y_{i},Y_{j})-\mu_{1101}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{11}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})D_{i}}{1-\pi(X_{i};\alpha)}\left\{h(Y_{j},Y_{i})-\mu_{1101}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1101}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1101}(X_{j},X_{i};\gamma)\bigg{]},$
for $1\leq i\neq j\leq n$. For $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$, it is straightforward to show that $2\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|\mathcal{O}_{i}\}=\phi_{h}^{s}(\mathcal{O}_{i})$ and $\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\allowbreak\mathcal{O}_{j};\theta^{*})\}=2^{-1}\mathbb{E}\{\phi_{h}^{s}(\mathcal{O}_{i})\}=\tau_{h,N}^{s}$. Thus, the numerators can be estimated using an empirical average of all pairs as $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})\equiv\binom{n}{2}^{-1}\sum_{i<j}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta})$ for $s\in\{10,00,11\}$. Therefore, we can obtain the following estimators for the PGCEs expressed in (13):
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 3.
Theorem 3 (Multiply robustness and efficiency under parametric working models)

### v4

@@ 3.2 Estimators based on nonparametric identification formulas
It is worth noting that our weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score for pairs, and can be viewed as an extension of the pairwise inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately; See the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
4 Multiply robust and efficient $U$-statistics
4.1 The efficient influence function for the PGCE estimand
@@
We first develop more efficient estimators with $\sqrt{n}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification. Second, the lower bound of the asymptotic variance is the variance of the EIF, providing a benchmark to define efficient estimators. Third, the EIF provides a device for incorporating data-adaptive machine learners for the nuisance parameters, suggesting feasible estimators that achieve the asymptotic variance lower bound with $\sqrt{n}$ convergence rate (Chernozhukov et al., 2018). We explore these properties by first formalizing the EIF for each PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$. We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following Theorem 1 is obtained.
@@
Under Assumptions 1 - 3, the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
@@
$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
@@
The EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1. The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al. (2022) with respect to the distribution of observed data, $\mathcal{O}$. Based on Theorems 1 and S3, the asymptotic variances of estimators for the PGCEs are lower bounded by the respective semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$ for $s\in\{10,00,11\}$.
@@ 4.2 Multiply robust estimators
The EIFs imply the semiparametric efficiency lower bound and pave the way for formulating optimal estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and require care for constructing the final estimator. We use $U$-statistics in constructing the final estimator, which could be shown to have the same influence function given in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that $\tau_{h}^{10}$ can be rewritten as
$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}}.$ (12)
[…]
@@
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg],$
[…]
$\widehat{\tau}_{h,\mathrm{tr}}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (13)
which is essentially a $U$-statistic but in the form of ratios. Estimators $\widehat{\tau}_{h,\mathrm{tr}}^{00}$ and $\widehat{\tau}_{h,\mathrm{tr}}^{11}$ are similarly constructed and defined in Section S4.1 of the Online Supplement. The asymptotic properties of these estimators are summarized in Theorem 2.
Theorem 2 (Multiply robustness and efficiency)